OpenStudy (beccab003):

Adding and subtracting radicals help! Medal to best answer. Thanks! I'll post example question below.

OpenStudy (beccab003):

\[\sqrt{98}-\sqrt{72}+\sqrt{32}\]

OpenStudy (anonymous):

9.89-8.4+5.6=7.09 first you have to find the square root of all numbers then you have subtract 8.4 from 9.89 then add 5.6 to that answer and you get 7.09

OpenStudy (beccab003):

Aren't you supposed to find the perfect square of the numbers? I have no idea where you got all those decimals.

OpenStudy (anonymous):

I have a weird calculator so I entered in your problem and that's what it gave me

OpenStudy (beccab003):

You can't get a correct answer by entering this problem into a calculator. It doesn't work like that because your supposed to add and simplify. You have to find the perfect square of the number because we need to only work with whole numbers. I just have a really hard time understanding how to find the perfect square.

OpenStudy (triciaal):

An easier approach, without electronic calculator what are factors of 98? what are factors of 72? what are factors of 32?